A copper wire has a resistance of 5 Ω at 20°C. What is its resistance if the temperature. Changed to 75°C


1
Expert's answer
2021-03-24T20:03:50-0400

Given,

Resistance of the copper wire "(R_o)=5\Omega"

"T_1= 20^\circ C"

"R_2=?"

"T_2=75^\circ C"

"R=R_o(1+\alpha dT)"

temperature coefficient for copper "(\alpha)= +0.39\times 10^{-2}\Omega/^\circ C"


Now, "R_2=5(1+0.39\times 10^{-2}\times (75-20))"


"\Rightarrow R_2=5(1+0.39\times 10^{-2}\times 55)\Omega"


"\Rightarrow R_2 =5\times 1.21\Omega"

"\Rightarrow R_2 = 6.05 \Omega"